# Semi-supervised and unsupervised learning for health indicator extraction from guided waves in aerospace composite structures

### v1

@@ Abstract
Health indicators (HIs) are central to diagnosing and prognosing the condition of aerospace composite structures, enabling efficient maintenance and operational safety. However, extracting reliable HIs remains challenging due to variability in material properties, stochastic damage evolution, and diverse damage modes. Manufacturing defects (e.g., disbonds) and in‑service incidents (e.g., bird strikes) further complicate this process. This study presents a comprehensive data‑driven framework that learns HIs via two learning approaches integrated with multi‑domain signal processing. Because ground‑truth HIs are unavailable, a semi-supervised and an unsupervised approach are proposed: (i) a diversity deep semi-supervised anomaly detection (Diversity-DeepSAD) approach augmented with continuous auxiliary labels used as hypothetical damage proxies, which overcomes the limitation of prior binary labels that only distinguish healthy and failed states while neglecting intermediate degradation, and (ii) a degradation-trend-constrained variational autoencoder (DTC-VAE), in which the monotonicity criterion is embedded via an explicit trend constraint. Guided waves with multiple excitation frequencies are used to monitor single‑stiffener composite structures under fatigue loading. Time, frequency, and time–frequency representations are explored, and per‑frequency HIs are fused via unsupervised ensemble learning to mitigate frequency dependence and reduce variance. Using fast Fourier transform features, the augmented Diversity-DeepSAD model achieved 81.6% performance, while DTC-VAE delivered the most consistent HIs with 92.3% performance, outperforming existing baselines. This work contributes: (i) a Diversity-DeepSAD variant for intermediate degradation modeling via continuous labels, (ii) the first application of Diversity-DeepSAD and DTC-VAE to guided wave monitoring of aerospace composites, and (iii) a framework for history-free HI estimation. These advancements provide reliable HIs, reducing maintenance costs and enhancing safety in aerospace structures.
@@
List of Symbols $c$ Hypersphere centre $D$ Depth of network $F_{all}$ Fitness function $F_{test}$ Test fitness function $f$ Frequency $H$ Number of statistical features $h$ Statistical feature $i,j$ Timestep $k$ Prognostic criteria weighting $\mathcal{L}$ Loss function $M$ Total number of samples MoMo Monotonicity $m$ Sample number NN Number of measurements (timesteps) $N_{f}$ Number of frequencies $n$ Number of labeled samples $P_{\theta}$ Decoder probability distribution PrPr Prognosability $Q_{\phi}$ Encoder probability distribution $r$ Rate of degradation $S$ Frequency domain statistical feature $s(f)$ Frequency domain signal $t_{i}$ Time at timestep $i$ TrTr Trendability $u$ Number of unlabeled samples $W_{l}$ Network weights of layer $l$ $X$ Time domain statistical feature $x(i)$ Time domain signal $x_{i}^{m,f}$ Features at timestep $i$ of specimen $m$ with measurement frequency $f$ $\hat{x}$ Reconstructed features $y_{i}^{m,f}$ Health indicator at timestep $i$ of specimen $m$ with measurement frequency $f$ $\tilde{y}$ Training label $\bar{y}$ Mean health indicator across folds $z$ Variational autoencoder latent variables $\alpha$ Kullback-Leibler divergence weighting $\beta$ Reconstruction loss weighting $\gamma$ Monotonicity constraint loss weighting $\epsilon$ Gaussian noise $\zeta_{k}$ $k$th eigenvalue of Gram matrix $\eta$ Labelled parameter loss weighting $\theta$ Neural network function $\lambda$ Diversity loss weighting $\mu$ Mean $\nu$ L2 regularisation weighting $\sigma$ Standard deviation $T$ Test specimen $\tau$ Training specimen $\phi$ Encoder parameters $\psi$ Decoder parameters $\omega_{f}$ Weight for frequency $f$
@@ 1 Introduction
The health of aerospace structures must be continuously monitored to ensure their safety and reliability. An effective monitoring and diagnostic system should detect damage early enough and issue warnings before it escalates into a dangerous situation [1]. Even earlier diagnostics using damage trend prediction, known as prognostics, can save significant time and cost by enabling timely repairs before a structure reaches an irreparable state [2, 3], in addition to improving reliability [4, 5, 6]. In this context, developing a health indicator (HI) is essential for both diagnostics and prognostics [7, 8, 9, 10]. A HI, being a value which quantifies the health of a structure, serves as a critical bridge between these two aspects towards condition-based maintenance (CBM) [11, 12]. However, deriving a truly comprehensive and accurate HI is exceptionally challenging for most objects, especially engineering systems, due to their inherent complexity [13]. The comprehensiveness of a HI refers to its ability to account for all potential types of damage that could affect a system’s health [14, 15, 13]. Tracking and quantifying all forms of degradation, even in simple objects made of isotropic materials, is often infeasible. Many degradation processes occur internally, beyond the reach of direct measurement. Despite the fact that sensors may monitor the effects of different types of damage on the sensed signals, meticulously identifying the damage remains challenging due to numerous known and unknown factors, which depend on the object under monitoring, type of sensors, and structural health monitoring (SHM) techniques employed. This complexity is further exacerbated in the case of advanced systems and materials [14, 16]. In summary, claiming to have completely true HI values for any object, especially a complex one like an engineering system, is an overstatement [13, 17].
@@
While a perfectly accurate HI is impractical, certain principles about HIs hold universally true. For example, the end of life (EoL) represents the complete loss of health in any system. Similarly, a system’s health inevitably decreases over time unless maintenance or self-healing mechanisms are applied. These principles form the foundation for evaluating a HI using key metrics. Prognostability (Pr), for instance, ensures that the failure state at EoL is represented by the same HI value across different systems [18, 19]. Monotonicity (Mo) emphasises the consistency of health degradation over time [20, 18, 21, 22]. Another desirable metric is trendability (Tr), which assesses whether the degradation trends of similar systems are consistent. Although uncertainties in manufacturing, loading conditions, and environmental factors may introduce variability, the general trends for similar systems should not diverge significantly. Therefore, ideal HIs for a group of similar systems should satisfy Mo, Pr, and Tr [19, 13, 23, 24] (considering all in a fitness score), with the first two being fact-based and the third being aspirational but still valuable.
@@
As already mentioned, developing or identifying HIs is a challenging task due to the inherent complexity and uncertainty of material degradation processes. This difficulty is amplified in anisotropic materials like composites [25] which are widely used in aerospace applications thanks to their high strength-to-weight ratio and ability to withstand directional stresses [26]. Unlike isotropic materials, composite materials exhibit more complex failure modes arising from their anisotropic properties and the potential interactions between different types of damage [27, 28, 15]. These unique characteristics make designing effective HIs for composite structures particularly intricate [14, 23]. Additionally, uncertainties arising during manufacturing (e.g., disbond defects) or operation (e.g., bird strikes) exacerbate this complexity [29, 30]. To enable in situ monitoring and account for these uncertainties, SHM techniques are essential. An effective SHM technique must be capable of covering most structural areas, particularly critical regions. Among these techniques, guided wave (GW)-based SHM stands out as a practical approach, where multiple sensors are strategically attached to critical regions to monitor the structure’s health. GWs are ultrasonic waves bounded by a structure’s material [30] that produce data from which the location and severity of damage to a structure can be determined [27, 31, 32, 33]. This enables SHM in otherwise inaccessible regions of complex composite structures [34]. As GW and most other SHM techniques rely on a network of sensors, model-based approaches face limitations in processing the large volume of complex, high-dimensional sensed signals. Furthermore, noise, multimodality, and the influence of environmental factors make it more challenging to generate HIs from raw GW data [15, 35, 36]. Consequently, data-driven and artificial intelligence (AI) approaches offer greater potential for handling such intricate datasets. Due to the unknown true health state of any engineering system, supervised models face limitations in constructing HIs. To address this challenge, semi-supervised or unsupervised AI models can be utilised for data fusion, enabling the generation of effective HIs [37, 38].
Semi-supervised models begin with concepts developed from supervised methods. For example, the support vector machine (SVM), introduced by Cortes and Vapnik [39], is a fundamental supervised classification algorithm. However, its binary classification ability falls short of generating meaningful HIs, as it ignores the complex gradient between healthy and failure states. Progress toward HI generation came with support vector data description (SVDD) [40] and its deep learning extension, Deep SVDD, proposed by Kim et al. in 2015 [41], which captures more intricate behaviors. For SHM, a semi-supervised model is preferred to effectively handle unlabeled data. Deep semi-supervised anomaly detection (DeepSAD), proposed by Ruff et al. in 2019 [42], extends Deep SVDD by incorporating unlabeled data points. While DeepSAD has proven effective in various non-SHM contexts, its application to HI generation poses challenges. For instance, Han et al. [43] used DeepSAD to detect anomalies in vehicle emissions, by setting a threshold anomaly score. They evaluated the performance of models by the area under the receiver operating characteristic curve, thus without considering the distribution of scores on either side of the threshold. This approach, although beneficial for anomaly detection, is inadequate for HI extraction due to the complete neglect of prognostic criteria. Similarly, DeepSAD has been applied with consideration only to the classification of anomalies in hydraulic systems [44]. Other approaches have sought to enhance semi-supervised anomaly detection. Gao et al. [45] developed ConNet, a robust deep anomaly detection model for sparsely labelled data. However, ConNet only considers positive, anomalous, labeled samples, which is unsuitable for HI generation as the healthy state is not considered.
Recent advancements have specifically targeted HI generation. Frusque et al. [46] applied DeepSAD to HI generation for rotating machinery, and it was observed that the objective function considers only the norm of the embedding, therefore producing low-rank results with some dimensions not fully utilised. To diversify the representation of degradation trends, they incorporated a diversity term in the loss function, creating Diversity-DeepSAD. For the PHME2010 milling dataset, Diversity-DeepSAD achieved promising results, with performances of 99% for Mo, 99% for Tr, and 94% for Pr, indicating room for improvement in Pr. However, when applied to thermal spray coating monitoring data, the model’s performance was lower, with 91% for Mo and 72% for Pr, while Tr was not reported. Based on the displayed HIs, the Tr score appears low. Notably, the coefficient of variation (CV), useful for comparing variability across datasets, was 24% for the thermal spray coating dataset. In contrast, aeronautical structures are significantly more challenging, with CV values of 51% for GW data from five single-stiffener composite panels [15] and 87% for acoustic emission data from 12 panels [23, 24]. Despite its potential, Diversity-DeepSAD has not yet been applied to aeronautical structures, which present unique challenges for HI construction and remaining useful life (RUL) prediction.
Moradi [13] developed various semi-supervised frameworks for single-stiffener composite structures. One approach utilised a multi-layer long short-term memory (LSTM) network, employing an intrinsically semi-supervised inductive learning technique, trained on acoustic emission-based time and frequency domain features [23]. This method achieved a fitness score of 93% under a rigorous leave-one-out cross-validation (LOOCV) process. However, the fine-tuning process to optimise the number of epochs should be done using only the validation specimen (unit), and another unit, different from the training/validation units, should be investigated to confirm the generalisability of this fine-tuned hyperparameter—something that was not considered in their study. To address this, Moradi et al. [24] introduced CEEMDAN-driven semi-supervised ensemble deep learning (CEEMDAN-SSEDL), which combined complete ensemble empirical mode decomposition with adaptive noise (CEEMDAN) for feature extraction and ensemble learning techniques to reduce randomness. By incorporating bidirectional LSTM (BiLSTM) layers during ensemble learning, this approach achieved a fitness score of 91.3%, while ensuring separate test units were considered. Moradi et al.’s [23] semi-supervised inductive learning approach was later applied by Frusque et al. [46] to the PHME2010 milling dataset, achieving a fitness score of 96%. However, the details of the architecture and the hyperparameter fine-tuning process were not reported. Recognising the limitations of historical-dependent models, Moradi et al. [37, 13] noted that such models often perform poorly in the absence of historical data. To overcome this, they introduced the Hilbert transform semi-supervised convolutional neural network (HT-SSCNN), designed to generate HIs from historical-independent GW data. Leveraging the GW monitoring technique, this approach achieved a 93% fitness score. Inspired by this, the methods in this paper are also designed to be historical-independent, relying solely on data from the current state of the structure. Given the demonstrated capabilities of anomaly detection algorithms, this study seeks to extend the application of Diversity-DeepSAD to the SHM of aerospace structures.
Previous applications of DeepSAD, however, have been limited to binary auxiliary labels distinguishing only healthy and failed states. This binary formulation neglects the intermediate degradation states that make up the majority of a structure’s lifetime, and in composites often leaves only a single label at failure, severely restricting the usefulness of labels. To address this gap, we propose the use of continuous auxiliary labels interpolated across the degradation process, providing richer supervision and smoother, more prognostic HIs.
Unsupervised models, by design, rely on feedback from the model output to learn without labeled data. A common technique is the autoencoder (AE), which uses a neural network to encode input data into a reduced-dimensional latent space and then reconstruct the input from this representation. Yang et al. [47] generated HIs for rotating machinery using sparse AEs, evaluating them based on the Mann-Kendall (MK) monotonicity metric. Lin and Tao [48] extended this approach by employing an ensemble of stacked AEs with linear targets, while Xu and Wang [49] enhanced stacked AEs with noise reduction using an exponential weight moving average model. In 2023, Xu et al. [50] implemented a deep convolutional AE and Mao et al. [51] applied tensor representation. These studies, however, focused on rotating machinery, where the vibrational behavior is more self-exposing and sensitive to damage. In contrast, aerospace structures do not exhibit such clear damage effects on the monitored signals, making the application of these methods more challenging. Variational autoencoder (VAE) architecture is similar to AE, but makes use of a probabilistic latent variable distribution in order to increase generability to data beyond the training set. In 2019, Ping et al. [52] utilised a VAE with a logarithmic normal distribution to generate HIs for turbofan engines (CMAPSS simulated dataset, which is simpler compared to experimental dataset [23, 53]). In 2020, Hemmer et al. [54] applied a VAE to rotating machinery at discrete damage levels. While generally increasing with damage level, the HIs do not conform to the prognostic criteria. Mitigating this, in 2022, Qin et al. [55] introduced a monotonic degradation constraint in the loss function, resulting in the degradation-trend-constrained VAE (DTC-VAE), which generates HIs that meet the Mo metric. Guo et al. [56] proposed a multiscale convolutional AE network for rotating machinery. Further studies in 2024 applied developments of VAE to the generation of HIs for bearings from vibrational data [57, 58, 59]. However, the application of VAEs for HI generation in aerospace structures remains unattempted. This paper seeks to address this gap by applying the DTC-VAE model to aerospace structures, aiming to achieve high HI evaluation criteria scores comparable to those achieved by semi-supervised methods. By leveraging the strengths of both Diversity-DeepSAD and DTC-VAE, the study aspires to advance HI generation for SHM in complex aeronautical environments.
@@
Furthermore, the calculation of the typical fitness score, which includes Mo, Pr, and Tr, is prone to bias toward the training units. This bias can lead to potential issues, as the model may produce highly correlated HIs during training, resulting in misleadingly high fitness scores. However, when faced with an unmatched HI from a specific unit during testing, the performance may fall short. To address this, the rectified evaluation criteria [15, 13] are adopted in this study, ensuring a more reliable assessment during the test phase. This adjustment provides a stronger foundation for comparison and enhances the practical reliability of the standard.
@@
To construct a historical-independent HI for aeronautical composite structures, GW-SHM emerges as a promising technique as mentioned earlier. However, each GW measurement requires data collection through a network of PZT sensors, resulting in an enormous volume of data to process. Feeding this data directly into deep learning models demands a high number of input layer nodes, which increases model complexity and computational requirements while reducing interpretability [36]. A more efficient alternative is to employ signal processing (SP) algorithms as a preprocessing step. SP algorithms often follow explicit solutions, producing globally applicable and faster outcomes. By extracting a set of statistical features from the processed signals, the size and complexity of the models can be significantly reduced. However, identifying which SP method yields the most effective features remains an open question. The proposed frameworks address this by first applying various SP algorithms to extract and compare features in terms of HIs criteria. The selected features are then fed into AI models to generate HIs, ensuring both efficiency and effectiveness in the process.
@@
    Introduction of two novel historical-independent HI models designed for reliable RUL estimation with reduced data inputs.
@@
    First application of DTC-VAE and Diversity-DeepSAD models to SHM of aerospace composite structures, specifically guided-wave SHM.
@@
    Novel use of intermediary proxy HIs as labels in a DeepSAD-based model, moving beyond binary healthy/failed labels, enabling richer semi-supervised HI construction.
@@
    Extension of historical-independent HI frameworks to semi- and unsupervised learning models.
@@
The methodology employed in this study is described in detail in Section 2, with the corresponding results presented in Section 3. A discussion of these findings is provided in Section 4, culminating in the conclusions outlined in Section 5.

### v2

@@ Abstract
Health indicators (HIs) are central to diagnosing and prognosing the condition of aerospace composite structures, enabling efficient maintenance and operational safety. However, extracting reliable HIs remains challenging due to variability in material properties, stochastic damage evolution, and diverse damage modes. Manufacturing defects (e.g. disbonds) and in‑service incidents (e.g. bird strikes) further complicate this process. This study presents a comprehensive data‑driven framework that learns HIs via two learning approaches integrated with multi‑domain signal processing. Because ground‑truth HIs are unavailable, a semi-supervised and an unsupervised approach are proposed: (i) a diversity deep semi-supervised anomaly detection (Diversity-DeepSAD) approach augmented with continuous auxiliary labels used as hypothetical damage proxies, which overcomes the limitation of prior binary labels and enables modelling of intermediate degradation, and (ii) a degradation-trend-constrained variational autoencoder (DTC-VAE), in which the monotonicity criterion is embedded via an explicit trend constraint. Guided waves with multiple excitation frequencies are used to monitor single‑stiffener composite structures under fatigue loading. Time, frequency, and time–frequency representations are explored, and per‑frequency HIs are fused via unsupervised ensemble learning to mitigate frequency dependence and reduce variance. Using fast Fourier transform features, the models achieved fitness scores of 81.6% (Diversity-DeepSAD) and 92.3% (DTC-VAE), indicating improved monotonicity and consistency over existing baselines. The proposed history-independent framework, supported by prognostic metrics–guided Bayesian optimisation and excitation frequency-agnostic HI fusion, enables the estimation of more robust HIs for aeronautical composite structures.
@@
List of Symbols $c$ Hypersphere centre $F_{all}$ Fitness function $F_{test}$ Test fitness function $f$ Frequency $H$ Number of eigenvalues $h$ Eigenvalue number $i,j$ Timestep $k$ Prognostic criteria weighting $\mathcal{L}$ Loss function $M$ Total number of specimens MoMo Monotonicity $m$ specimen number NN Number of measurements (timesteps) $N_{f}$ Number of frequencies $n$ Number of labeled samples PrPr Prognosability $Q_{\phi}$ Encoder probability distribution $r$ Rate of degradation $S$ Frequency domain statistical feature $s(f)$ Frequency domain signal $t_{i}$ Time at timestep $i$ TrTr Trendability $u$ Number of unlabeled samples $X$ Time domain statistical feature $x(i)$ Time domain signal $x_{i}^{m,f}$ Features at timestep $i$ of specimen $m$ with measurement frequency $f$ $\hat{x}$ Reconstructed features $y_{i}^{m,f}$ Health indicator at timestep $i$ of specimen $m$ with measurement frequency $f$ $\tilde{y}$ Auxiliary label $\bar{y}$ Mean health indicator across folds $z$ Autoencoder latent variables $\alpha$ Kullback-Leibler divergence weighting $\beta$ Reconstruction loss weighting $\gamma$ Monotonicity constraint loss weighting $\epsilon$ Small number to prevent zero errors $\varepsilon$ Gaussian noise $\zeta_{h}$ $h^{th}$ eigenvalue of Gram matrix $\eta$ Labelled parameter loss weighting $\theta$ Neural network function $\lambda$ Diversity loss weighting $\mu$ Mean $\nu$ L2 regularisation weighting $\sigma$ Standard deviation $\phi$ Encoder parameters $\psi$ Decoder parameters $\omega_{f}$ Weight for frequency $f$
@@ 1 Introduction
The health of aerospace structures must be continuously monitored to ensure their safety and reliability. An effective monitoring and diagnostic system should detect damage early enough to issue warnings before it escalates into a dangerous situation [jimenez2020towards]. Even earlier diagnostics using damage trend prediction, known as prognostics, can save significant time and cost by enabling timely repairs before a structure reaches an irreparable state [sikorska2011prognostic, badihi2022comprehensive], in addition to improving reliability [jia2021sample, Guo2024, wu2025multi]. In this context, developing a health indicator (HI) is essential for both diagnostics and prognostics [qiu2003robust, guo2017recurrent, khan2023review, Wei2025]. An HI, being a value which quantifies the health of a structure, serves as a critical bridge between these two aspects towards condition-based maintenance (CBM) [zhang2022marine, huang2024prognostics]. However, deriving a truly comprehensive and accurate HI is exceptionally challenging for most objects, especially engineering systems, due to their inherent complexity [moradi2024designing]. The comprehensiveness of an HI refers to its ability to account for all potential types of damage that could affect a system’s health [moradi2024designing, beaumont2020structural, Moradi2024]. Tracking and quantifying all forms of degradation, even in simple objects made of isotropic materials, is often infeasible. Many degradation processes occur internally, beyond the reach of direct measurement. Although sensors may monitor the effects of different types of damage on the sensed signals, meticulously identifying the damage remains challenging due to numerous known and unknown factors, which depend on the object under monitoring, type of sensors, and structural health monitoring (SHM) techniques employed. This complexity is further exacerbated in the case of advanced systems and materials [beaumont2020structural, ferreira2022remaining]. In summary, claiming to have completely true HI values for any object, especially a complex one like an engineering system, is an overstatement [moradi2024designing, Huang2024].
@@
While a perfectly accurate HI is impractical, certain principles about HIs hold universally true. For example, the end of life (EoL) represents the complete loss of health in any system. Similarly, a system’s health inevitably decreases over time unless maintenance or self-healing mechanisms are applied. These principles form the foundation for evaluating an HI using key metrics. Prognostability ($Pr$), for instance, ensures that the failure state at EoL is represented by the same HI value across different systems [coble2009identifying, coble2010merging]. Monotonicity ($Mo$) emphasises the consistency of health degradation over time [saxena2008metrics, coble2009identifying, Gonzalez-Muniz2022, Li2023]. Another desirable metric is trendability ($Tr$), which assesses whether the degradation trends of similar systems are consistent. Although uncertainties in manufacturing, loading conditions, and environmental factors may introduce variability, the general trends for similar systems should not diverge significantly. Therefore, ideal HIs for a group of similar systems should satisfy $Mo$, PrPr, and $Tr$ [coble2010merging, moradi2024designing, Moradi2023Intelligent, moradi2025novel], all of which are considered in the fitness score.
@@
As already mentioned, developing or identifying HIs is a challenging task due to the inherent complexity and uncertainty of material degradation processes. This difficulty is amplified in anisotropic materials like composites [senthilkumar2021nondestructive] which are widely used in aerospace applications thanks to their high strength-to-weight ratio and ability to withstand directional stresses [hassani2021structural]. Unlike isotropic materials, composite materials exhibit more complex failure modes arising from their anisotropic properties and the potential interactions between different types of damage [deluca2018, Saeedifar2020, Moradi2024]. These unique characteristics make designing effective HIs for composite structures particularly intricate [beaumont2020structural, Moradi2023Intelligent]. Additionally, uncertainties arising during manufacturing (e.g. disbond defects) or operation (e.g. bird strikes) exacerbate this complexity [Dienel2019, Aujoux2023]. To enable in situ monitoring and account for these uncertainties, SHM techniques are essential. An effective SHM technique must be capable of covering most structural areas, particularly critical regions. Among these techniques, guided wave (GW)-based SHM stands out as a practical approach, where multiple sensors are strategically attached to critical regions to monitor the structure’s health. GWs are ultrasonic waves bounded by a structure’s material [Aujoux2023] that produce data from which the location and severity of damage to a structure can be determined [deluca2018, Flynn2011, Memmolo2018, Kralovec2020]. This enables SHM in otherwise inaccessible regions of complex composite structures [Janardhan2014]. As GW and most other SHM techniques rely on a network of sensors, model-based approaches face limitations in processing the large volume of complex, high-dimensional sensed signals. Furthermore, noise, multimodality, and the influence of environmental factors make it more challenging to generate HIs from raw GW data [Moradi2024, Yang2020, Yang2023]. Consequently, data-driven and artificial intelligence (AI) approaches offer greater potential for handling such intricate datasets. Due to the unknown true health state of any engineering system, supervised models face limitations in constructing HIs. To address this challenge, semi-supervised or unsupervised AI models can be utilised for data fusion, enabling the generation of effective HIs [Moradi2023SemiSupverised, Li2024].
@@
Furthermore, the calculation of the typical fitness score, which includes $Mo$, PrPr, and $Tr$, is prone to bias toward the training units. This bias can lead to potential issues, as the model may produce highly correlated HIs during training, resulting in misleadingly high fitness scores. However, when faced with an unmatched HI from a specific unit during testing, the performance may fall short. To address this, the rectified evaluation criteria [moradi2024designing, Moradi2024] are adopted in this study, ensuring a more reliable assessment during the test phase. This adjustment provides a stronger foundation for comparison and enhances the practical reliability of the standard.
@@
To construct a history-independent HI for aeronautical composite structures, GW-SHM emerges as a promising technique as mentioned earlier. However, each GW measurement requires data collection through a network of PZT sensors, resulting in an enormous volume of data to process. Feeding this data directly into deep learning models demands a high number of input layer nodes, which increases model complexity and computational requirements while reducing interpretability [Yang2023]. A more efficient alternative is to employ signal processing (SP) algorithms as a preprocessing step. SP algorithms often follow explicit solutions, producing globally applicable and faster outcomes. By extracting a set of statistical features from the processed signals, the size and complexity of the models can be significantly reduced. However, identifying which SP method yields the most effective features remains an open question. The proposed frameworks address this by first applying various SP algorithms to extract and compare features in terms of fitness criteria. The selected features are then fed into AI models to generate HIs, ensuring both efficiency and effectiveness in the process.
@@
    Extension of the Diversity-DeepSAD model into augmented Diversity-DeepSAD by embedding continuous degradation-progression labels instead of binary $\pm 1$ labels in its loss function, enabling modelling of intermediate fatigue states essential for composites.
@@
    Development of a degradation-trend-constrained variational autoencoder (DTC-VAE) architecture, in which a monotonicity constraint is embedded directly in the latent space to enforce consistent HI evolution, for guided wave monitoring of aerospace composites.
@@
    Use of Bayesian hyperparameter optimisation guided by a composite fitness score based on monotonicity, prognosability, and trendability, requiring construction of full run-to-failure HI trajectories for all training and validation units.
@@
    Implementation of a frequency-agnostic unsupervised HI fusion mechanism, weighted by composite fitness score, providing a principled alternative to selecting a single excitation frequency in guided wave SHM.
  5. (v)
    Comprehensive comparison of semi- and unsupervised learning models for generating history-independent HIs, highlighting their behaviour, limitations, and applicability across guided wave SHM scenarios.
@@
A review of fundamental literature follows in Section 2. The methodology employed in this study is described in detail in Section 3, with the corresponding results presented in Section 4. A discussion of these findings is provided in Section 5, culminating in the conclusions outlined in Section 6.
2 Literature study
In this section, we first review semi-supervised models relevant to HI generation, followed by a discussion of unsupervised approaches. Semi-supervised models begin with concepts developed from supervised methods. For example, the support vector machine (SVM), introduced by Cortes and Vapnik [cortes1995], is a fundamental supervised classification algorithm. However, its binary classification ability falls short of generating meaningful HIs, as it ignores the complex gradient between healthy and failure states. Progress toward HI generation came with support vector data description (SVDD) [tax2004] and its deep learning extension, Deep SVDD, proposed by Kim et al. in 2015 [kim2015], which captures more intricate behaviours. For SHM, a semi-supervised model is preferred to effectively handle unlabeled data. Deep semi-supervised anomaly detection (DeepSAD), proposed by Ruff et al. in 2019 [Ruff2019], extends Deep SVDD by incorporating unlabeled data points. While DeepSAD has proven effective in various non-SHM contexts, its application to HI generation poses challenges. For instance, Han et al. [Han2025] used DeepSAD to detect anomalies in vehicle emissions, by setting a threshold anomaly score. They evaluated the performance of models by the area under the receiver operating characteristic curve, thus without considering the distribution of scores on either side of the threshold. This approach, although beneficial for anomaly detection, is inadequate for HI extraction due to the complete neglect of prognostic criteria. Similarly, DeepSAD has been applied with consideration only to the classification of anomalies in hydraulic systems [dong2023]. Other approaches have sought to enhance semi-supervised anomaly detection. Gao et al. [Gao2021] developed ConNet, a robust deep anomaly detection model for sparsely labelled data. However, ConNet only considers positive, anomalous, labeled samples, which is unsuitable for HI generation as the healthy state is not considered.
Recent advancements have specifically targeted HI generation. Frusque et al. [Frusque2023] applied DeepSAD to HI generation for rotating machinery, and it was observed that the objective function considers only the norm of the embedding, therefore producing low-rank results with some dimensions not fully utilised. To diversify the representation of degradation trends, they incorporated a diversity term in the loss function, creating Diversity-DeepSAD. For the PHME2010 milling dataset, Diversity-DeepSAD achieved promising results, with performances of 99% for $Mo$, 99% for $Tr$, and 94% for $Pr$, indicating room for improvement in $Pr$. However, when applied to thermal spray coating monitoring data, the model’s performance was lower, with 91% for $Mo$ and 72% for $Pr$, while $Tr$ was not reported. Based on the displayed HIs, the $Tr$ score appears low. Notably, the coefficient of variation (CV), useful for comparing variability across datasets, was 24% for the thermal spray coating dataset. In contrast, aeronautical structures are significantly more challenging, with CV values of 51% for GW data from five single-stiffener composite panels [Moradi2024] and 87% for acoustic emission data from 12 specimens [Moradi2023Intelligent, moradi2025novel]. Despite its potential, Diversity-DeepSAD has not yet been applied to aeronautical structures, which present unique challenges for HI construction and remaining useful life (RUL) prediction.
Moradi [moradi2024designing] developed various semi-supervised frameworks for single-stiffener composite structures. One approach utilised a multi-layer long short-term memory (LSTM) network, employing an intrinsically semi-supervised inductive learning technique, trained on acoustic emission-based time and frequency domain features [Moradi2023Intelligent]. This method achieved a fitness score of 93% under a rigorous leave-one-out cross-validation (LOOCV) process. However, the fine-tuning process to optimise the number of epochs should be done using only the validation specimen (unit), and another unit, different from the training/validation units, should be investigated to confirm the generalisability of this fine-tuned hyperparameter—something that was not considered in their study. To address this, Moradi et al. [moradi2025novel] introduced CEEMDAN-driven semi-supervised ensemble deep learning (CEEMDAN-SSEDL), which combined complete ensemble empirical mode decomposition with adaptive noise (CEEMDAN) for feature extraction and ensemble learning techniques to reduce randomness. By incorporating bidirectional LSTM (BiLSTM) layers during ensemble learning, this approach achieved a fitness score of 91.3%, while ensuring separate test units were considered. Moradi et al.’s [Moradi2023Intelligent] semi-supervised inductive learning approach was later applied by Frusque et al. [Frusque2023] to the PHME2010 milling dataset, achieving a fitness score of 96%. However, the details of the architecture and the hyperparameter fine-tuning process were not reported. Recognising the limitations of history-dependent models, Moradi et al. [moradi2024designing, Moradi2023SemiSupverised] noted that such models often perform poorly in the absence of historical data. To overcome this, they introduced the Hilbert transform semi-supervised convolutional neural network (HT-SSCNN), designed to generate HIs from history-independent GW data. Leveraging the GW monitoring technique, this approach achieved a 93% fitness score. Inspired by this, the methods in this paper are also designed to be history-independent, relying solely on data from the current state of the structure. Given the demonstrated capabilities of anomaly detection algorithms, this study seeks to extend the application of Diversity-DeepSAD to the SHM of aerospace structures.
Previous applications of DeepSAD, however, have been limited to binary auxiliary labels distinguishing only healthy and failed states. This binary formulation neglects the intermediate degradation states that make up the majority of a structure’s lifetime, and in composites often leaves only a single label at failure, severely restricting the usefulness of labels. To address this gap, we introduce continuous auxiliary labels interpolated across the degradation process, applied only at early-life (e.g. 0–25%) and late-life (e.g. 75–100%) stages, to provide richer supervision and produce smoother HIs.
Unsupervised models, by design, rely on feedback from the model output to learn without labeled data. A common technique is the autoencoder (AE), which uses a neural network to encode input data into a reduced-dimensional latent space and then reconstruct the input from this representation. Yang et al. [yang2018] generated HIs for rotating machinery using sparse AEs, evaluating them based on the Mann-Kendall (MK) monotonicity metric. Lin and Tao [lin2019] extended this approach by employing an ensemble of stacked AEs with linear targets, while Xu and Wang [xu2022] enhanced stacked AEs with noise reduction using an exponential weight moving average model. In 2023, Xu et al. [xu2023hi] implemented a deep convolutional AE and Mao et al. [mao2023] applied tensor representation. These studies, however, focused on rotating machinery, where the vibrational behaviour is more self-exposing and sensitive to damage. In contrast, aerospace structures do not exhibit such clear damage effects on the monitored signals, making the application of these methods more challenging. Variational autoencoder (VAE) architecture is similar to AE, but makes use of a probabilistic latent variable distribution in order to increase generability to data beyond the training set. In 2019, Ping et al. [Ping2019] utilised a VAE with a logarithmic normal distribution to generate HIs for turbofan engines (CMAPSS simulated dataset, which is simpler compared to experimental dataset [Moradi2023Intelligent, moradi2025constructing]). In 2020, Hemmer et al. [hemmer2020] applied a VAE to rotating machinery at discrete damage levels. While generally increasing with damage level, the HIs do not conform to the prognostic criteria. Mitigating this, in 2022, Qin et al. [Qin2022] introduced a monotonic degradation constraint in the loss function, resulting in the degradation-trend-constrained VAE (DTC-VAE), which generates HIs that meet the $Mo$ metric. Guo et al. [guo2022] proposed a multiscale convolutional AE network for rotating machinery. Further studies in 2024 applied developments of VAE to the generation of HIs for bearings from vibrational data [li2024rotating, li2024vae, milani2024]. However, the application of VAEs for HI generation in aerospace structures remains unattempted. This paper seeks to address this gap by applying the DTC-VAE model to GW signals measured from aerospace structures, aiming to achieve high quality HIs. By leveraging the strengths of both Diversity-DeepSAD and DTC-VAE, the study aspires to advance HI generation for complex aeronautical structures.